OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

hero (hero):

Hint: Divide both sides by -2 to get -2(x-6) = x - 3

OpenStudy (anonymous):

how did you get this

hero (hero):

I divided both sides by -2

hero (hero):

-2/-2 = 1 4/-2 = -2

hero (hero):

Now distribute -2 across x - 6 to get: -2x + 12 = x - 3

hero (hero):

Afterwards, add 2x to both sides and 3 to both sides to get: 15 = 3x Finish the rest